OpenStudy (anonymous):

Find the quotient (1/5) / (3/4)

OpenStudy (anonymous):

1/5 = 0.2

OpenStudy (anonymous):

3/4 = 0.75

OpenStudy (igreen):

The answer is \(\Large\frac{4}{15}\).

OpenStudy (anonymous):

thanks !